OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function. f(x) = -2x^4 + 4x^3 + 3x^2 + 18

OpenStudy (anonymous):

Lol do you need help with this one ?

OpenStudy (anonymous):

please !!!!!!!!!

OpenStudy (anonymous):

what are the factors of the leading coefficient? what are the factors of the constant term?

OpenStudy (anonymous):

Similar problem as an example ; f(x) = -2x^3 + 3x^2 -2x +8 Solution: Factors of the constant term (8) / Factors of the leading coefficient (-2) so that would be: 1, 2, 4, 8 / 1, 2 So then each numerator over each denominator is a rational zero.... +/- 1/1, +/- 2/1, +/- 4/1, +/- 8/1, +/- 1/2, +/- 2/2, +/- 4/2, and +/- 8/2 Simplified: +/- 1, +/- 2, +/- 4, +/- 8, +/- 1/2, +/-1, +/- 2, and +/- 4 There are some duplicates so you can "drop" them... So your final answer is: +/- 1/2, +/- 1, +/- 2, +/- 4, +/- 8

OpenStudy (anonymous):

omg this is so confusing

OpenStudy (anonymous):

f(x) = -2x^4 + 4x^3 + 3x^2 + 18 constant term? Factors of the leading coefficient?

OpenStudy (anonymous):

factors of 2: 1,2 factors of 18: 1,2,3,4,6,18 so some examples include 1/1,1/2,2/1,2/2,3/1,3/2,...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what are constant term? Factors of the leading coefficient? and do i need to do that for every number?

OpenStudy (anonymous):

@jenny1003

OpenStudy (anonymous):

-2 and 18 Sorry I'm helping alot of ppl at the same time

OpenStudy (anonymous):

factors of 2: 1,2 factors of 18: 1,2,3,4,6,18 factors of 4: 1,2,4 factors of 13: 1,3,

OpenStudy (anonymous):

3 not 13

OpenStudy (anonymous):

The leading term of a polynomial is the term of highest degree. The leading coefficient of a polynomial is the coefficient of the leading term. The degree of a polynomial is the degree of the leading term.

OpenStudy (anonymous):

leading term 4 leading coefficient -2

OpenStudy (anonymous):

I'm lost right now

OpenStudy (anonymous):

i am too

OpenStudy (anonymous):

@200205650 Please help....Srry

OpenStudy (anonymous):

So you are pretty much solving for x in this equation right?

OpenStudy (anonymous):

i don't really know :O

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

This video might help in solving this problem. And yes you are trying to find what are all the numbers that can equal x. Factoring is one way of solving this problem. I can help you if you don't know how to factor this problem.

OpenStudy (anonymous):

@ganeshie8 @nincompoop

OpenStudy (anonymous):

@Peaches15

OpenStudy (anonymous):

@DanTheMan99

OpenStudy (anonymous):

i just saw the video, but still i can't understand

OpenStudy (anonymous):

Did you watch the whole thing?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

i didn't

OpenStudy (anonymous):

i saw the beginning

OpenStudy (anonymous):

Okay why don't you watch that video and I'll try to solve this problem using factoring. Is that fine?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

@DanTheMan99 check this out http://www.wolframalpha.com/input/?i=f%28x%29+%3D+-2x%5E4+%2B+4x%5E3+%2B+3x%5E2+%2B+18

OpenStudy (anonymous):

\[-2x^4+4x^3+3x^2+18\]\[-2x^3(x-2)+3(x^2+6)\]\[(-2x^3+3)(x-2)(x^2+6)\] Equation #1:\[(-2x^3+3)=0\]\[-2x^3=-3\]\[x^3=\frac{ 3 }{ 2 }\]\[x=0.72112\] Equation#2: \[(x-2)=0\]\[x=2\] Equation #3:\[(x^2+6)=0\]\[x^2=-6\]\[x=\sqrt{6}i\]

OpenStudy (anonymous):

thank you guys so much

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Greatly done, @200205650

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

+ 1/2, + 1, + 2, + 4, + 8, - 1/2, - 1, - 2, - 4, and -8.

OpenStudy (anonymous):

i don't know how to answer

OpenStudy (anonymous):

like how to put it into words

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

@nono266 please help me

OpenStudy (anonymous):

what should i write

oregonduck (oregonduck):

you have it written fine

OpenStudy (anonymous):

i know but whats my explanation

OpenStudy (anonymous):

like what proof can i give

oregonduck (oregonduck):

there is no need to do so

OpenStudy (anonymous):

are you sure

oregonduck (oregonduck):

100% positive

oregonduck (oregonduck):

is this flvs?

OpenStudy (anonymous):

yeah

oregonduck (oregonduck):

ok then yeah you do not need to

oregonduck (oregonduck):

for sure

oregonduck (oregonduck):

it is fine if you have the right answer then they know you understand it

OpenStudy (anonymous):

ok thanks

oregonduck (oregonduck):

do i get 10 owl bucks because i need them

oregonduck (oregonduck):

?????????????

oregonduck (oregonduck):

do you have paypal?

OpenStudy (anonymous):

no

oregonduck (oregonduck):

can you ask the question for me and i go under it. ask a qualified helper question

oregonduck (oregonduck):

ask the question as ALGEBRA 1 HELP THEN I WILL ADD THE REST